OpenStudy (anonymous):

Question 1 If we toss 3 coins, what is the probability that no tails will appear? Answer 7/8 3/4 1/8 1/2 Question 2 If we toss 3 coins, what is the probability that exactly 1 tail will appear? Answer 3/8 1/2 3/4 1/8 Question 3 If we toss a pair of dice, what is the probability that a sum of 7 will appear? Answer 1/36 7/36 1/2 1/6

OpenStudy (anonymous):

@ganeshie8

OpenStudy (anonymous):

Q1 1-(0.5)^3 = 7/8 Q2 I'm not sure but I'm pretty certain that it's 1/8 Q3 6/36 = 1/6

OpenStudy (anonymous):

are you sure the other answers are correct?

OpenStudy (anonymous):

I am fairly certain but it's always good to make sure with others

OpenStudy (anonymous):

Concentrationalizing are you going to reply to my question please do :]

OpenStudy (anonymous):

When we have only a few amount of outcomes, we can count them out pretty easily. Let's look at all of the possibilities that arise from flipping 3 coins. H = heads, T = tails: HHH TTT HHT TTH HTH THT THH HTT So we have 8 possibilities, so all of our probabilities are out of 8. We see that 7 of those 8 have at least 1 tails, so the probability of NO tails is 7/8 If we now look at, out of those 3 options, how many times a tails appears once, we can count and see it occurs 3 times, so we have 3/8 for question 2. As for the dice roll, we can look at this drawing:  Okay, quite a bit going on, but I have every possibility, all 36, for a roll of two die. On the top running right and down are the sums of dice rolls on the diagonals I circled and on the left going down I have the probabilities of the dice rolls on the circled diagonals. As you can see, I have a diagonal circled that has a number 7 next to it. If you follow that same diagonal down, you can see that the diagonal has the probability of 6/36 = 1/6